OpenStudy (anonymous):

A cylinder has a surface area of 402 cm^2. The height is three times greater than the radius. What is the height of the cylinder? 8 cm, 10.48 cm, 12 cm, or 16.97 cm

OpenStudy (anonymous):

r = radius 3r = height (h) SA = 2(pi)(r)(h) + 2(pi)(r^2) SA = 2(pi)(r)(3r) + 2(pi)(r^2) SA = 6(pi)(r^2) + 2(pi)(r^2) SA = 8(pi)r^2 402 = 8(pi)r^2 402 = 8(3.14)r^2 402 = 25.12r^2 16.00318 = r^2 rounding 16.00318 = 16 16 = r^2 r = 4

OpenStudy (gudden):

formula for surface area of cylinder is 2 pi r^2 + 2 pi r h where r = radius and h = height and let radius be x and than according to given height = 3x now try equating these.. and try the question yourself

OpenStudy (anonymous):

britney explained it but it was a little hard to follow... can someone explain it a little more clearly?